OpenStudy (anonymous):

Rationalize the denominator of square root of negative 9 over open parentheses 4 minus 7 I close parentheses minus open parentheses 6 minus 6 I close parentheses

OpenStudy (mertsj):

\[\frac{\sqrt{-9}}{(4-7i)-(6-6i)}=\frac{3i}{-2-i}\]

OpenStudy (mertsj):

\[\frac{3i}{-2-i}\times\frac{-2+i}{-2+i}=\frac{-6i+3i^2}{4+1}=\frac{-3-6i}{5}\]